OpenStudy (anonymous):

find the sum of 2sqrt6 and sqrt24

OpenStudy (anonymous):

\[2\sqrt{6}+\sqrt{24}=2\sqrt{6}+2\sqrt{6}=4\sqrt{6}\]

OpenStudy (anonymous):

and gimme a medal :D

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

thanx :D

OpenStudy (experimentx):

likes the second post ... LOL

OpenStudy (anonymous):

HAHA